OpenStudy (anonymous):

How would i write f(x)=2x+30 in inverse. the answer is f(x)=1/2(x-30) but i dont know the steps to get that answer

OpenStudy (anonymous):

this is the steps---while f(x) = y f(x) = 2x + 30 y = 2x + 30 y - 30 = 2x (y - 30)1/2 = x substitute the value of y for x. therefore (x - 30)1/2 = f(x) inverse